An aqueous solution of sulfuric acids, H2SO4, is reacted with with a 0.122 M solution of sodium hydroxide, NaOH.

a. Write the balance equation.

b. How many ml of sodium hydroxide are needed to neutralize 27.0 ml of 0.080 M of the sulfuric acid?
Answer #1

a)
Balanced chemical equation is:
H2SO4 + 2 NaOH ---> Na2SO4 + 2 H2O

b)
Here:
M(H2SO4)=0.08 M
M(NaOH)=0.122 M
V(H2SO4)=27.0 mL

According to balanced reaction:
2*number of mol of H2SO4 =1*number of mol of NaOH
2*M(H2SO4)*V(H2SO4) =1*M(NaOH)*V(NaOH)
2*0.08 M *27.0 mL = 1*0.122M *V(NaOH)
V(NaOH) = 35.4 mL

Answer: 35.4 mL